OpenStudy (anonymous):

last month it was estimated that a lake contained 3500 rainbow trout. over a three day period a park ranger caught, tagged and realeased 100 fish. then after two weeks for random mixing she caught 100 more rainbow trout & found three of them had tags. a) what's the probability of catching a tagged fish? b) the answer is: you have to assume that the population is 3500, it remians stable & the fish are well mixed. c) based on the number of tagged fish she caught two weeks later whats the park rangers experimental probability? anyone help with a and c pleeeease?<3

OpenStudy (anonymous):

Single Event Probability Formula : Probability of event A that occurs P(A) = n(A) / n(S). Probability of event A that does not occur P(A') = 1 - P(A). Multiple Event Probability Formula : Probability of event A that occurs P(A) = n(A) / n(S). Probability of event A that does not occur P(A') = 1 - P(A). Probability of event B that occurs P(B) = n(B) / n(S). Probability of event B that does not occur P(B') = 1 - P(B). Probability that both the events occur P(A ∩ B) = P(A) x P(B). Probability that either of event occurs P(A ∪ B) = P(A) + P(B) - P(A ∩ B). Conditional Probability P(A | B) = P(A ∩ B) / P(B). where, n(A) - number of occurrence in Event A, n(B) - number of occurrence in Event B, n(S) - total number of possible outcomes.

OpenStudy (anonymous):

whaaaaaaaaaat?!

OpenStudy (anonymous):

i think reading the last 3 lines should help, then go to the top

OpenStudy (anonymous):

miss taylor..... Probability of event A occuring = n(A)/n(S) n(S) is the total no of outcomes, or in this case the total no of fish..which is 3500 n(a) is the number of tagged fish..no of favorable outcomes..which is 100

OpenStudy (anonymous):

so the theoretical probability is 100/3500 = 1/ 35 = 2.86 %

OpenStudy (anonymous):

thanks him, i gotta go cause ive been here about 6hrs too tired to sit in front of a computer all day sseyall

OpenStudy (anonymous):

thnx...noprobs